# On the Consistency of Multilingual Context Utilization in Retrieval-Augmented Generation

### v3

@@ 4.2 Interpretability-based Assessment
To further verify our findings that the passage language is not a barrier to LLMs’ understanding capability of the multilingual retrieved passages, we adopt MIRAGE (Qi et al., 2024), a model internal-based method for attributing model responses to the retrieved passages in RAG systems. Generally, it consists of two components: (1) CTI for detecting contextual sensitivity for the generated sentence and (2) CCI for attributing the detected sentences back to each retrieved passage. Given the single-passage setup, in this section we only use the CTI module for evaluating the passage dependency of the model response. For each generated token, this module measures the shift in output probability distribution when no context vs. one passage is provided, measured by KL divergence (Kullback and Leibler, 1951), while keeping the generated sentence prefix fixed. If at least one token is higher than an empirically set CTI threshold, the generated sentence is marked as sensitive to the context provided in the prompt.
@@ 5.1 Accuracy Results
Table 3 presents the results, including the no-context baseline and single in-/out-language gold passage results as computed in Section 4, to enable comparison (see Appendix I for full language-specific results). For this analysis, we also consider as valid the questions that were answered correctly but in the wrong language, as they also reflect a proper understanding of the context by the model. Interestingly, models provided with 3 out-language gold passages achieve higher accuracy than when provided with a single in-language gold passage in the query language, emphasizing the potential of cross-lingual retrieval for mRAG. As expected, the presence of distractors leads to lower accuracy. Notably, this is true for all models, datasets, and setups. However, the effect is considerably stronger in MKQA than in GMMLU-Choice, likely due to the stricter lexical-matching metric adopted for MKQA. We also verify that a higher proportion of distractors (3/4 vs. 1/4) is much more harmful for answer accuracy, which confirms the importance of having access to a high-quality cross-lingual retriever (Chirkova et al., 2024). When comparing the drop between in-language distractors and out-language distractors, we find that in-language distractors have a larger impact in most cases, matching our hypothesis that this is a particularly challenging scenario for LLMs. However, differences are small in many cases, indicating the language of the distractor is not a major issue for multi-passage mRAG.
@@ 5.2 Interpretability-based Evaluation
We adopt once again MIRAGE (Qi et al., 2024) to understand how the internal model dynamics are affected by our various simulated multi-passage mRAG scenarios. We sample 50 instances from each dataset and use MIRAGE to attribute Aya-expanse-8B responses to the provided passages via contrastive feature attribution (Yin and Neubig, 2022). Then, we compute # Contextual: the average number of distracting passages that contain at least one contextual cue for the produced answer (i.e. a token marked by CCI in MIRAGE), and # Influential: the average number of distractors that receive a higher sum of CCI attribution scores than all gold passages for each query.
@@ 6 Conclusion
At the same time, we also detected some important directions for future improvement. Firstly, poor generation abilities in many languages push the models to respond in a different language than that of the query, resulting in answers that would be deemed useless by most end-users. Importantly, we showed that this also happens when the retrieval works optimally. This suggests that, rather than just trying to optimize the retriever, it may be more effective to invest on the model generation abilities in a specific (set of) user language(s) –for instance by continued pre-training (Fujii et al., 2024; Gao et al., 2024) on generic corpora of those languages– or to apply techniques that push the model to decode in a given language, such as contrastive decoding (Li et al., 2023; O’Brien and Lewis, 2023). Secondly, the presence of distracting passages (i.e., relevant to the query topic, but not directly functional to answer it) in the context can have a very negative effect on answer accuracy in open-domain QA. While this effect is rather similar regardless of the distractors’ language, it does highlight the importance of carefully ranking the retrieved passages and to aim for precision when selecting which passages are provided to the model.
@@ Appendix E Prompts and Instructions
To ensure the model responses are always in the query language, we follow previous works (Chirkova et al., 2024; Zhang et al., 2024) and adopt language-specific instructions to explicitly and implicitly guide the model to generate responses in the user-readable language. The examples in English, Spanish, and Chinese are listed in Table 9 and Table 10.
@@ Appendix F Error Analysis on XQUAD
While our MKQA and GMMLU results strongly suggest our studied LLMs can understand the provided passages regardless of their language, the in-/out-language gap in XQUAD remains unexplained. To address this, we conduct a manual error analysis on XQUAD with Aya-Expanse-8B, focusing on a random sample of 20 Spanish and 20 Chinese queries that were answered correctly when provided with in-language passages, but wrongly with out-language passages. In most cases, we observe that models successfully understood the context and generated a proper response, however, this response did not perfectly match the gold answer provided in the dataset. This can be due to the presence of synonyms or slight paraphrases of the gold answer, or –in the case of languages with different scripts– to transliteration variations (Knight and Graehl, 1997). For instance, the gold answer for a Spanish question is ‘evolución de la lengua y la literatura alemanas’ (i.e. ‘evolution of the German language and literature’). In the in-language setup, the model manages to generate this exact string as it is included in the provided Spanish passage. However, when the same passage is provided in English, the model generates the semantically equivalent phrase ‘… evolución del idioma y la literatura alemana…’, or ‘…desarrollo del idioma y la literatura alemana…’ when the passage is provided in Chinese. Similarly, for a Chinese query with gold answer ‘亚里士多德宇宙学’ (i.e. ‘Aristotelian cosmology’), model responses slightly differ when provided with different out-language passages (e.g. ‘亚里士多德宇宙论’, ‘阿里斯托的宇宙论’, or ‘阿里斯托特利宇宙论’ with English, Arabic, or Greek passage respectively), all of which are correct translations of ‘Aristotelian cosmology’. While this issue can always affect lexical-matching evaluation, it is particularly severe in XQUAD as many answers in this dataset are named entities or sentence segments due to the extractive nature of the task, which in turn causes an underestimation of the models capability.
@@ Appendix G Advanced System Prompting
In our main experiments, we follow the previous works (Chirkova et al., 2024; Zhang et al., 2024) and adopt the direct prompt. To test if a stronger prompt could mitigate language-mismatch errors, we add a two-step instruction that first allows the model to answer in any appropriate language, then explicitly translates the answer into the query language. Formally: ‘Write a high-quality answer to the given question using the provided search results. Please respond in English. Specifically, please follow the two steps below. Step 1: Generate a complete answer to the question in any appropriate language. Step 2: Translate your entire answer into clear, natural‐sounding English.’

### v4

@@ 4.2 Interpretability-based Assessment
To further verify our findings that the passage language is not a barrier to LLMs’ understanding capability of the multilingual retrieved passages, we adopt MIRAGE (qi-etal-2024-model), a model internal-based method for attributing model responses to the retrieved passages in RAG systems. Generally, it consists of two components: (1) CTI for detecting contextual sensitivity for the generated sentence and (2) CCI for attributing the detected sentences back to each retrieved passage. Given the single-passage setup, in this section we only use the CTI module for evaluating the passage dependency of the model response. For each generated token, this module measures the shift in output probability distribution when no context vs. one passage is provided, measured by KL divergence (kldiv), while keeping the generated sentence prefix fixed. If at least one token is higher than an empirically set CTI threshold, the generated sentence is marked as sensitive to the context provided in the prompt.
@@ 5.1 Accuracy Results
Table 3 presents the results, including the no-context baseline and single in-/out-language gold passage results as computed in Section 4, to enable comparison (see Appendix I for full language-specific results). For this analysis, we also consider as valid the questions that were answered correctly but in the wrong language, as they also reflect a proper understanding of the context by the model. Interestingly, models provided with 3 out-language gold passages achieve higher accuracy than when provided with a single in-language gold passage in the query language, emphasizing the potential of cross-lingual retrieval for mRAG. As expected, the presence of distractors leads to lower accuracy. Notably, this is true for all models, datasets, and setups. However, the effect is considerably stronger in MKQA than in GMMLU-Choice, likely due to the stricter lexical-matching metric adopted for MKQA. We also verify that a higher proportion of distractors (3/4 vs. 1/4) is much more harmful for answer accuracy, which confirms the importance of having access to a high-quality cross-lingual retriever (chirkova-etal-2024-retrieval). When comparing the drop between in-language distractors and out-language distractors, we find that in-language distractors have a larger impact in most cases, matching our hypothesis that this is a particularly challenging scenario for LLMs. However, differences are small in many cases, indicating the language of the distractor is not a major issue for multi-passage mRAG.
@@ 5.2 Interpretability-based Evaluation
We adopt once again MIRAGE (qi-etal-2024-model) to understand how the internal model dynamics are affected by our various simulated multi-passage mRAG scenarios. We sample 50 instances from each dataset and use MIRAGE to attribute Aya-expanse-8B responses to the provided passages via contrastive feature attribution (yin-neubig-2022-interpreting). Then, we compute # Contextual: the average number of distracting passages that contain at least one contextual cue for the produced answer (i.e. a token marked by CCI in MIRAGE), and # Influential: the average number of distractors that receive a higher sum of CCI attribution scores than all gold passages for each query.
@@ 6 Conclusion
At the same time, we also detected some important directions for future improvement. Firstly, poor generation abilities in many languages push the models to respond in a different language than that of the query, resulting in answers that would be deemed useless by most end-users. Importantly, we showed that this also happens when the retrieval works optimally. This suggests that, rather than just trying to optimize the retriever, it may be more effective to invest on the model generation abilities in a specific (set of) user language(s) –for instance by continued pre-training (fujii2024continual; gao2024multilingual) on generic corpora of those languages– or to apply techniques that push the model to decode in a given language, such as contrastive decoding (li-etal-2023-contrastive; o2023contrastive). Secondly, the presence of distracting passages (i.e., relevant to the query topic, but not directly functional to answer it) in the context can have a very negative effect on answer accuracy in open-domain QA. While this effect is rather similar regardless of the distractors’ language, it does highlight the importance of carefully ranking the retrieved passages and to aim for precision when selecting which passages are provided to the model.
@@ Appendix E Prompts and Instructions
To ensure the model responses are always in the query language, we follow previous works (chirkova-etal-2024-retrieval; zhang-etal-2024-respond) and adopt language-specific instructions to explicitly and implicitly guide the model to generate responses in the user-readable language. The examples in English, Spanish, and Chinese are listed in Table 9 and Table 10.
@@ Appendix F Error Analysis on XQUAD
While our MKQA and GMMLU results strongly suggest our studied LLMs can understand the provided passages regardless of their language, the in-/out-language gap in XQUAD remains unexplained. To address this, we conduct a manual error analysis on XQUAD with Aya-Expanse-8B, focusing on a random sample of 20 Spanish and 20 Chinese queries that were answered correctly when provided with in-language passages, but wrongly with out-language passages. In most cases, we observe that models successfully understood the context and generated a proper response, however, this response did not perfectly match the gold answer provided in the dataset. This can be due to the presence of synonyms or slight paraphrases of the gold answer, or –in the case of languages with different scripts– to transliteration variations (knight-graehl-1997-machine). For instance, the gold answer for a Spanish question is ‘evolución de la lengua y la literatura alemanas’ (i.e. ‘evolution of the German language and literature’). In the in-language setup, the model manages to generate this exact string as it is included in the provided Spanish passage. However, when the same passage is provided in English, the model generates the semantically equivalent phrase ‘… evolución del idioma y la literatura alemana…’, or ‘…desarrollo del idioma y la literatura alemana…’ when the passage is provided in Chinese. Similarly, for a Chinese query with gold answer ‘亚里士多德宇宙学’ (i.e. ‘Aristotelian cosmology’), model responses slightly differ when provided with different out-language passages (e.g. ‘亚里士多德宇宙论’, ‘阿里斯托的宇宙论’, or ‘阿里斯托特利宇宙论’ with English, Arabic, or Greek passage respectively), all of which are correct translations of ‘Aristotelian cosmology’. While this issue can always affect lexical-matching evaluation, it is particularly severe in XQUAD as many answers in this dataset are named entities or sentence segments due to the extractive nature of the task, which in turn causes an underestimation of the models capability.
@@ Appendix G Advanced System Prompting
In our main experiments, we follow the previous works (chirkova-etal-2024-retrieval; zhang-etal-2024-respond) and adopt the direct prompt. To test if a stronger prompt could mitigate language-mismatch errors, we add a two-step instruction that first allows the model to answer in any appropriate language, then explicitly translates the answer into the query language. Formally: ‘Write a high-quality answer to the given question using the provided search results. Please respond in English. Specifically, please follow the two steps below. Step 1: Generate a complete answer to the question in any appropriate language. Step 2: Translate your entire answer into clear, natural‐sounding English.’